OpenStudy (algtrigcalc):

Please help! An operation * has been defined over the set of natural numbers. (3*2)*2; x*y = (x+y)/(xy) How do I determine whether this is commutative/associative or not?

zepdrix (zepdrix):

What is the (3*2)*2 in front for?

OpenStudy (algtrigcalc):

It's find (3*2)*2

zepdrix (zepdrix):

\[\large\rm \color{orangered}{(3*2)}*2\quad=\quad\color{orangered}{\frac{3+2}{3\cdot2}}*2\]Were you able to figure that part of the problem out?

OpenStudy (algtrigcalc):

Yes. So we substitute the answer into the equation again since its *2 right?

zepdrix (zepdrix):

Yes, first simplify the orange part though. I guess that works out to 5/6 right? Oh wait wait wait... this operation is defined over the `natural numbers`... not necessarily the rational numbers (which 5/6 belongs to). Hmm

zepdrix (zepdrix):

Ya I don't think this operation is defined for (3*2)*2 Hmmmm D:

OpenStudy (algtrigcalc):

oh ok. Can you tell me in general how to check if something is commutative or associative when an operation is defined?

OpenStudy (algtrigcalc):

so for commutative, would it be (y+x)/(xy) = (x+y)/(xy)

zepdrix (zepdrix):

Let \(\rm x,y,z\in\mathbb N\) Then, \(\large\rm (x*y)*z=?\) Work that out and determine whether or not it is equal to \(\large\rm x*(y*z)\). That's what we need for `associativity`, ya?

zepdrix (zepdrix):

Yes, you have the right idea. Let \(\rm x,y\in\mathbb N\) Then determine whether or not \(\large\rm x*y=y*x\)

OpenStudy (algtrigcalc):

How do I know what * is?

zepdrix (zepdrix):

Well it was defined above, yes?\[\large\rm x*y=\frac{x+y}{xy}\]then just apply a simple step here (since we know that our normal addition and multiplication are commutative,\[\large\rm \large\rm x*y=\frac{x+y}{xy}=\frac{y+x}{yx}=y*x\]

OpenStudy (algtrigcalc):

Ohh I see. ((x+y)/(xy)) * 2 = (x/(xy)+((y/xy) * 2) And find if that is equal for it to be associative?

zepdrix (zepdrix):

No no, let that last value be some unknown from the naturals. Otherwise we know that associativity works with 2, but not necessarily anything else. And also, I don't think you want to break up the fractions.. seems like that's going to make it really really complicated. I mean, for one thing, we don't know if * is distributive.

zepdrix (zepdrix):

\[\large\rm (x*y)*z=\frac{x+y}{xy}*z=\frac{\frac{x+y}{xy}+z}{\left(\frac{x+y}{xy}\right)z}\]Get something like that, ya?

zepdrix (zepdrix):

Do the same for this \(\large\rm x*(y*z)\). Start in the brackets.

OpenStudy (algtrigcalc):

Sorry, I'm not sure if I understand....how do I write (y*z) if I can't break up the operation?

zepdrix (zepdrix):

We defined * above, yes? :o\[\large\rm (y*z)=\frac{y+z}{yz}\]You having trouble with this weirdly defined operation?

zepdrix (zepdrix):

For this problem, Star just means, add in the numerator, multiply in the denominator.

OpenStudy (algtrigcalc):

Ohh, ok. so \[x*(y*z) = x* \frac{ y+z}{ yz } = \frac{ \frac{ y+z }{ yz }+z }{ \frac{ y+z }{y z } *z }\]

zepdrix (zepdrix):

What...? What happened to your x? :(

zepdrix (zepdrix):

And why is the fraction on the left side?

OpenStudy (algtrigcalc):

Sorry, I put in the wrong key. the z is suppose to be the x

zepdrix (zepdrix):

I'm not sure why you put the x's on the right side... that's rather confusing. And also, do not use * for multiplication, that's extra confusing :D lol

zepdrix (zepdrix):

\[\large\rm x*(y*z)=x*\frac{y+z}{yz}=\frac{x+\frac{y+z}{yz}}{x\cdot\frac{y+z}{yz}}\]ok good.\[\large\rm (x*y)*z=\frac{x+y}{xy}*z=\frac{\frac{x+y}{xy}+z}{\left(\frac{x+y}{xy}\right)z}\]We need to determine whether or not these big fractions that we ended up with are equivalent. So I guess we need to simplify them further :[

zepdrix (zepdrix):

Err maybe we don't need to. I think it's pretty clear from this point that they are not going to be equal. Umm

OpenStudy (algtrigcalc):

Thank you so much so staying on this problem so long to help me! So we should keep it in variables to find if the operation is commutative or associative rather than substituting numbers in?

zepdrix (zepdrix):

Yes. If you plug in specific numbers, you're able to determine that you have associativity *sometimes*, but not necessarily all the time. You may have just gotten lucky and found a few numbers that are associative. Leave them unknown so they can represent any natural from the set. So what did we end up with here? I think we determined that * is `associative` but not `commutative`? I hope we did everything alright. :P

zepdrix (zepdrix):

Is this like.. Abstract Algebra or something? Maybe ask someone or teacher about the (3*2)*2 thing. This isn't my strongest subject so it'd be good to get another opinion :D

OpenStudy (algtrigcalc):

It's part of the section addition and multiplication properties of real numbers. Did you mean to say that it is commutative but it's not associative?

zepdrix (zepdrix):

Ahh yes, good call.